A certain compound has the molecular formula X4O6 .If 10gm of compound contains 5.62gm of X, the atomic mass of X is....?

Grade:11

1 Answers

Arun
3 years ago
Mass of the Compound (X₄O₆) = 10 g.
Mass of the Element X in the Compound = 6.06 g.
 
Let the Atomic Mass of the Element X be a.
∴ Mass of X in the compound = 4a.
 
∴ Molecular Mass of the Compound = a × 4 + 16 × 6
   = 4a + 96
  
 
Now,
Mass of X/Mass of the compound = At. Mass of the X/Mol. Mass of the Compound.
⇒ 6.06/10 = 4a/(4a + 96)
⇒ 6.06/10 = 4a/4(a + 24)
⇒ 0.606 = a/(a + 24)
⇒ 0.606(a + 24) = a
⇒ 0.606a + 14.544 = a
⇒ a - 0.606a = 14.544
⇒ 0.394a = 14.544
⇒ a = 14.544/0.394
∴ a = 36.9
∴ a ≈ 37 a.m.u.
 
 
Hence, the Atomic Mass of the Element X is 37 a.m.u.